Sketch and angle theater in standing position such that data has the least possible positive measure in the 5, -12 is on the terminal side of theta. Then find the exact values of the Six trigonometric functions for theta.

first graph and then find the answers for the six trigonometric functions.
Answer #1

Date: Date: 11 (5,-22) in Peat standard position 5, 125 M D OAB op? - ABZ OB? A2 = 52 + 122 OA2= 25+ 198 - on = 169=13 OF 20-